# FaithfulRAG: Fact-Level Conflict Modeling for Context-Faithful Retrieval-Augmented Generation

### v1

@@ Abstract
Large language models (LLMs) augmented with retrieval systems have demonstrated significant potential in handling knowledge-intensive tasks. However, these models often struggle with unfaithfulness issues, generating outputs that either ignore the retrieved context or inconsistently blend it with the LLM’s parametric knowledge. This issue is particularly severe in cases of knowledge conflict, where the retrieved context conflicts with the model‘s parametric knowledge. While existing faithful RAG approaches enforce strict context adherence through well-designed prompts or modified decoding strategies, our analysis reveals a critical limitation: they achieve faithfulness by forcibly suppressing the model’s parametric knowledge, which undermines the model‘s internal knowledge structure and increases the risk of misinterpreting the context. To this end, this paper proposes FaithfulRAG, a novel framework that resolves knowledge conflicts by explicitly modeling discrepancies between the model‘s parametric knowledge and retrieved context. Specifically, FaithfulRAG identifies conflicting knowledge at the fact level and designs a self-thinking process, allowing LLMs to reason about and integrate conflicting facts before generating responses. Extensive experiments demonstrate that our method outperforms state-of-the-art methods. The code is available at https://github.com/DeepLearnXMU/Faithful-RAG.
@@
Qinggang Zhang1thanks: Equal contribution., Zhishang Xiang211footnotemark: 1, Yilin Xiao3, Le Wang4, Junhui Li5, Xinrun Wang6, Jinsong Su1,7thanks: Corresponding author. 1School of Informatics, Xiamen University, China 2Institute of Artificial Intelligence, Xiamen University, China 3The Hong Kong Polytechnic University, China  4Migu Meland Co., Ltd  5Soochow University, China 6Singapore Management University, Singapore  7Shanghai Artificial Intelligence Laboratory {zqg.zhang, xzs.xiang}@hotmail.com jssu@xmu.edu.cn
@@ 1 Introduction
However, we identified a critical limitation in existing faithful methods through a thorough analysis: while they can achieve context faithfulness, they often do so at the cost of increased risk of misinterpreting the context. As evidenced by our experimental results in Figure 2, state-of-the-art context-faithful methods reduce unfaithful errors by 6.65% but simultaneously cause a 6.42% rise in incorrect match errors on average. This occurs because existing methods Zhou et al. (2023); Ying et al. (2024) attempt to achieve faithfulness by forcibly suppressing the model’s confidence in its parametric knowledge without properly understanding and analyzing the differences between contextual information and the model’s inherent knowledge. Such suppression compromises the model‘s ability to critically evaluate and reconcile discrepancies, leading to degraded comprehension, logical incoherence, and a higher likelihood of aligning with incorrect contextual information.
@@ 2 Problem Statement
In our work, we focus on the generation stage and guide LLM to generate more faithful responses by accurately interpreting the context with the model‘s parametric knowledge.
@@ 3 Preliminary Study
Case 2: Incorrect-match Error. The LLM updates its parametric knowledge but incorrectly learns from the misleading context. To explore the underlying reasons, we evaluate state-of-the-art faithful RAG methods on the MuSiQue and SQuAD benchmarks Ying et al. (2024) and present the error distribution in Figure 2. The results reveal a critical limitation of existing models that they improve faithfulness at the cost of increasing the risk of misinterpreting the context. Specifically, we have the following observations:
Obs. 1. The vanilla RAG model (origin) exhibits a high proportion of Case 1 errors, with 13.6% on MuSiQue and 16.4% on SQuAD, while Case 2 errors remain relatively low at 5.2% and 7.1%, respectively. This suggests that the model primarily suffers from LLMs‘ inherent bias toward their parametric knowledge. Such a tendency aligns with prior findings Xie et al. (2024) that LLMs tend to prioritize pre-trained knowledge over contextual evidence when faced with conflicting information.
@@ Discussion.
Existing faithful RAG models Zhou et al. (2023); Yuan et al. (2024) reduce over-confidence errors but simultaneously cause a significant rise in incorrect match errors. This occurs because existing methods attempt to achieve faithfulness by forcibly suppressing the model’s confidence in its parametric knowledge without properly understanding and analyzing the discrepancies between context and the model’s inherent knowledge. Such suppression weakens the model‘s ability to critically evaluate interactions between parametric knowledge and contextual evidence, leading to either rigid adherence to outdated parametric knowledge or severe contextual overfitting, where the model passively adopts retrieved claims without critically evaluating their correctness.
@@ 4.1 Self-Fact Mining
It surfaces the LLM‘s implicit logical structure by identifying required knowledge domains and their interdependencies. Specifically, we prompt the LLM to identify the conceptual and factual prerequisites for answering $Q$, yielding a set of abstract, high-level knowledge:
@@ 4.2 Contextual Knowledge Alignment
To resolve knowledge conflicts while preserving logical coherence, in this section, FaithfulRAG aligns the model‘s externalized self-facts with the retrieved context through a structured, interpretable process. The alignment is structured as follows:
@@ 5.1 Experiment Settings
Evaluation Metrics and Implementation Details: Following previous studies, we evaluate all models using accuracy (ACC), where a model‘s response is considered correct only if it contains the ground truth answer. For the MuSiQue and SQuAD, we add $M_{R}$ (Memorization Ratio) Longpre et al. (2021) to measure context faithfulness. To ensure reproducibility, all models were evaluated using deterministic decoding (temperature $=$ 0). Further metrics and details are in the Appendix C.3.
@@ 5.4 Alation Study (Q3)
To evaluate the contributions of FaithfulRAG‘s core components, we systematically ablate three key components: (i) Self-Fact Mining, (ii) Self-Think module, and (iii) Chain-of-Thought (CoT), generating 7 variants as shown in Table 3. Each variant was tested under knowledge-conflict scenarios to isolate its impact on performance. We have the following findings.
Obs. 6. The ablation of Self-Knowledge Extraction and Self-Context Generation reveals that both modules are critical for precise knowledge alignment. Removing Self-Knowledge Extraction degrades the model‘s ability to analyze questions comprehensively, leading to a 1.1% average accuracy drop, as the LLM fails to identify relevant parametric facts. Conversely, removing Self-Context Generation—which converts abstract self-knowledge into actionable context—causes a larger 1.9% accuracy decline, demonstrating that raw parametric claims lack utility without contextual grounding.
Obs. 7. Ablating the full Self-Think module results in a 22.2% average accuracy drop, as simply prepending self-aligned context to the original context fails to resolve conflicts dynamically. Replacing Think stage with Special Annotation (explicitly marking key facts) reduces accuracy by 6.7%, proving that passive highlighting cannot replicate active reasoning. Similarly, substituting structured reasoning with naive Chain-of-Thought (CoT) decreases accuracy by 5.9%, confirming that explicit guidance (e.g., conflict anticipation steps) is essential. These results validate that Self-Think‘s diagnostic workflow (not just attention mechanisms or generic reasoning) drives reliable conflict resolution.
Obs. 8. While CoT enhances reasoning in some domains, naive application under knowledge conflicts causes a 17.1% accuracy drop, as the LLM grows distrustful of conflicting contexts. FaithfulRAG addresses this by integrating CoT with conflict anticipation: the LLM first identifies potential discrepancies in self-aligned contexts before reasoning, mitigating distrust. Notably, even without CoT, FaithfulRAG‘s accuracy decreases by only 2.1%, demonstrating its robustness. This highlights that CoT must be tailored to handle knowledge conflicts, and FaithfulRAG‘s structured reasoning framework achieves this adaptation.
To summarize, the combined ablation results reveal that FaithfulRAG‘s components operate synergistically. Self-Fact Mining provides the foundation for precise fact alignment, while Self-Think enables dynamic conflict resolution. Without either, the model reverts to the limitations of suppression-based methods (e.g., over-reliance on context or parametric knowledge). This interdependence validates the framework‘s design hypothesis: diagnostic reconciliation of conflicts instead of suppression is key to balancing faithfulness and accuracy.
@@ 6 Related Work
Context-Faithful Method. Some methods enhance context faithfulness through fine-tuning. KAFT Li et al. (2023) fine-tune small parameter models like T5 using counterfactual contexts, SFR-RAG  Nguyen et al. (2024)is a 9B models trained with various RAG-domain datasets. However, such fine-tuning approaches require substantial computational resources and are often difficult to transfer to other models. Among non-fine-tuning approaches, one category focuses on carefully designed prompt templates. For example, One method reconstruct the context as the narrator’s statement and then inquire about the narrator‘s viewpoint, encouraging the model to seek external perspectives Zhou et al. (2023). Another method employs "Role Play" interventions to alter LLMs’ decision-making styles when facing knowledge conflicts. Ying et al. (2024) However, these methods heavily rely on the model‘s inherent reasoning abilities and often suffer from limited generalizability. Another category of non-fine-tuning approaches modifies the model‘s decoding strategy to improve its reliance on context. CADShi et al. (2023)contrasts output probabilities with and without context to enhance contrastive decoding, amplifying the contextual probability distribution without accounting for conflicting contexts. In contrast, COIECDYuan et al. (2024) detects knowledge conflicts by measuring entropy changes during generation and dynamically adjusts the decoding strategy accordingly.
@@ 7 Conclusions
While FaithfulRAG advances error mitigation in text-based retrieval-augmented generation, its current scope is limited to textual inputs and does not yet support multimodal information (e.g., images, audio, or structured data). Extending the framework to incorporate multimodal input would enable a more comprehensive assessment of how models navigate and resolve conflicts between heterogeneous knowledge sources. Given that real-world information is often multimodal, such an extension could improve FaithfulRAG‘s ability to detect and reconcile inconsistencies that arise from modality-specific biases or divergent contextual signals.
In addition, multimodal integration would enhance the applicability of the framework to domains where information synthesis from multiple sources is crucial, such as medical diagnosis (where textual reports, imaging data and patient history must be jointly considered), autonomous systems (which rely on the integration of visual, auditory, and textual signals for decision making), and interactive AI (where understanding user intent often involves processing speech, gestures, and textual input). Addressing these challenges would not only improve FaithfulRAG‘s robustness but also contribute to broader advancements in multimodal retrieval-augmented generation, cross-modal reasoning, and trustworthy AI.
@@ A.2 What are advantages of FaithfulRAG?
Robustness in Non-Conflict Scenarios. A good faithful model should improve performance in knowledge-conflict scenarios while maintaining desirable performance in non-conflict scenarios. While most existing methods suffer in non-conflict settings (e.g., KRE‘s 13.8% drop on SQuAD-golden), our FaithfulRAG achieves the best performance on both SQuAD-golden and MuSiQue-golden. It gets 96.6% on SQuAD-golden (+1.4% over origin) and 85.3 on MuSiQue-golden (+1.3%), proving its robustness and capability to avoid overfitting to either LLM’s parametric knowledge or incorrect contextual information.
@@
Despite recent advances in retrieval mechanisms, a persistent challenge remains: a significant gap exists between the quality of the retrieved context and the factual accuracy of the generated output. While the context retrieved by the model may be highly relevant, the generative models often struggle to faithfully incorporate this context, resulting in generated text that may diverge from factual correctness. Recent empirical evidence reveals the critical gap between retrieval quality and generation faithfulness. The generation stage determines how this context is interpreted and integrated with the model‘s parametric knowledge. Existing methods often fail at this stage due to (i) Over-confidence: many RAG systems rigidly prioritize the retrieved context by suppressing parametric knowledge, leading to blind acceptance of flawed or conflicting information. (ii) Logical incoherence: Without critical reasoning, models struggle to resolve contradictions, resulting in inconsistent or unsafe outputs. By targeting generation, in this paper, we address the root cause of unfaithfulness and guide the LLM to dynamically reconcile knowledge conflicts.
@@ C.4 Implementation Details of Ablation Study
Finally, we conduct an ablation study on Reasoning stage. We substituted the structured reasoning with a naive Chain-of-Thought (CoT) approach, which does not provide explicit structured guidance for the LLM‘s reasoning. This modification makes the model rely solely on its implicit inference capabilities. The prompt is as follows:

### v2

@@ Abstract
Large language models (LLMs) augmented with retrieval systems have demonstrated significant potential in handling knowledge-intensive tasks. However, these models often struggle with unfaithfulness issues, generating outputs that either ignore the retrieved context or inconsistently blend it with the LLM’s parametric knowledge. This issue is particularly severe in cases of knowledge conflict, where the retrieved context conflicts with the model’s parametric knowledge. While existing faithful RAG approaches enforce strict context adherence through well-designed prompts or modified decoding strategies, our analysis reveals a critical limitation: they achieve faithfulness by forcibly suppressing the model’s parametric knowledge, which undermines the model’s internal knowledge structure and increases the risk of misinterpreting the context. To this end, this paper proposes FaithfulRAG, a novel framework that resolves knowledge conflicts by explicitly modeling discrepancies between the model’s parametric knowledge and retrieved context. Specifically, FaithfulRAG identifies conflicting knowledge at the fact level and designs a self-thinking process, allowing LLMs to reason about and integrate conflicting facts before generating responses. Extensive experiments demonstrate that our method outperforms state-of-the-art methods. The code is available at https://github.com/DeepLearnXMU/Faithful-RAG.
@@
Qinggang Zhang1thanks: Equal contribution., Zhishang Xiang211footnotemark: 1, Yilin Xiao3, Le Wang4, Junhui Li5, Xinrun Wang6, Jinsong Su1,7thanks: Corresponding author. 1School of Informatics, Xiamen University 2Institute of Artificial Intelligence, Xiamen University 3The Hong Kong Polytechnic University  4Migu Meland Co., Ltd  5Soochow University 6Singapore Management University  7Shanghai Artificial Intelligence Laboratory {zqg.zhang, xzs.xiang}@hotmail.com jssu@xmu.edu.cn
@@ 1 Introduction
However, we identified a critical limitation in existing faithful methods through a thorough analysis: while they can achieve context faithfulness, they often do so at the cost of increased risk of misinterpreting the context. As evidenced by our experimental results in Figure 2, state-of-the-art context-faithful methods reduce unfaithful errors by 6.65% but simultaneously cause a 6.42% rise in incorrect match errors on average. This occurs because existing methods Zhou et al. (2023); Ying et al. (2024) attempt to achieve faithfulness by forcibly suppressing the model’s confidence in its parametric knowledge without properly understanding and analyzing the differences between contextual information and the model’s inherent knowledge. Such suppression compromises the model’s ability to critically evaluate and reconcile discrepancies, leading to degraded comprehension, logical incoherence, and a higher likelihood of aligning with incorrect contextual information.
@@ 2 Problem Statement
In our work, we focus on the generation stage and guide LLM to generate more faithful responses by accurately interpreting the context with the model’s parametric knowledge.
@@ 3 Preliminary Study
Case 2: Incorrect-match Error. The LLM updates its parametric knowledge but incorrectly learns from the misleading context.
To explore the underlying reasons, we evaluate state-of-the-art faithful RAG methods on the MuSiQue and SQuAD benchmarks Ying et al. (2024) and present the error distribution in Figure 2. The results reveal a critical limitation of existing models that they improve faithfulness at the cost of increasing the risk of misinterpreting the context. Specifically, we have the following observations:
Obs. 1. The vanilla RAG model (origin) exhibits a high proportion of Case 1 errors, with 13.6% on MuSiQue and 16.4% on SQuAD, while Case 2 errors remain relatively low at 5.2% and 7.1%, respectively. This suggests that the model primarily suffers from LLMs’ inherent bias toward their parametric knowledge. Such a tendency aligns with prior findings Xie et al. (2024) that LLMs tend to prioritize pre-trained knowledge over contextual evidence when faced with conflicting information.
@@ Discussion.
Existing faithful RAG models Zhou et al. (2023); Yuan et al. (2024) reduce over-confidence errors but simultaneously cause a significant rise in incorrect match errors. This occurs because existing methods attempt to achieve faithfulness by forcibly suppressing the model’s confidence in its parametric knowledge without properly understanding and analyzing the discrepancies between context and the model’s inherent knowledge. Such suppression weakens the model’s ability to critically evaluate interactions between parametric knowledge and contextual evidence, leading to either rigid adherence to outdated parametric knowledge or severe contextual overfitting, where the model passively adopts retrieved claims without critically evaluating their correctness.
@@ 4.1 Self-Fact Mining
It surfaces the LLM’s implicit logical structure by identifying required knowledge domains and their interdependencies. Specifically, we prompt the LLM to identify the conceptual and factual prerequisites for answering $Q$, yielding a set of abstract, high-level knowledge:
@@ 4.2 Contextual Knowledge Alignment
To resolve knowledge conflicts while preserving logical coherence, in this section, FaithfulRAG aligns the model’s externalized self-facts with the retrieved context through a structured, interpretable process. The alignment is structured as follows:
@@ 5.1 Experiment Settings
Evaluation Metrics and Implementation Details: Following previous studies, we evaluate all models using accuracy (ACC), where a model’s response is considered correct only if it contains the ground truth answer. For the MuSiQue and SQuAD, we add $M_{R}$ (Memorization Ratio) Longpre et al. (2021) to measure context faithfulness. To ensure reproducibility, all models were evaluated using deterministic decoding (temperature $=$ 0). Further metrics and details are in the Appendix C.3.
@@ 5.4 Alation Study (Q3)
To evaluate the contributions of FaithfulRAG’s core components, we systematically ablate three key components: (i) Self-Fact Mining, (ii) Self-Think module, and (iii) Chain-of-Thought (CoT), generating 7 variants as shown in Table 3. Each variant was tested under knowledge-conflict scenarios to isolate its impact on performance. We have the following findings.
Obs. 6. The ablation of Self-Knowledge Extraction and Self-Context Generation reveals that both modules are critical for precise knowledge alignment. Removing Self-Knowledge Extraction degrades the model’s ability to analyze questions comprehensively, leading to a 1.1% average accuracy drop, as the LLM fails to identify relevant parametric facts. Conversely, removing Self-Context Generation—which converts abstract self-knowledge into actionable context—causes a larger 1.9% accuracy decline, demonstrating that raw parametric claims lack utility without contextual grounding.
Obs. 7. Ablating the full Self-Think module results in a 22.2% average accuracy drop, as simply prepending self-aligned context to the original context fails to resolve conflicts dynamically. Replacing Think stage with Special Annotation (explicitly marking key facts) reduces accuracy by 6.7%, proving that passive highlighting cannot replicate active reasoning. Similarly, substituting structured reasoning with naive Chain-of-Thought (CoT) decreases accuracy by 5.9%, confirming that explicit guidance (e.g., conflict anticipation steps) is essential. These results validate that Self-Think’s diagnostic workflow (not just attention mechanisms or generic reasoning) drives reliable conflict resolution.
Obs. 8. While CoT enhances reasoning in some domains, naive application under knowledge conflicts causes a 17.1% accuracy drop, as the LLM grows distrustful of conflicting contexts. FaithfulRAG addresses this by integrating CoT with conflict anticipation: the LLM first identifies potential discrepancies in self-aligned contexts before reasoning, mitigating distrust. Notably, even without CoT, FaithfulRAG’s accuracy decreases by only 2.1%, demonstrating its robustness. This highlights that CoT must be tailored to handle knowledge conflicts, and FaithfulRAG’s structured reasoning framework achieves this adaptation.
To summarize, the combined ablation results reveal that FaithfulRAG’s components operate synergistically. Self-Fact Mining provides the foundation for precise fact alignment, while Self-Think enables dynamic conflict resolution. Without either, the model reverts to the limitations of suppression-based methods (e.g., over-reliance on context or parametric knowledge). This interdependence validates the framework’s design hypothesis: diagnostic reconciliation of conflicts instead of suppression is key to balancing faithfulness and accuracy.
@@ 6 Related Work
Context-Faithful Method. Some methods enhance context faithfulness through fine-tuning. KAFT Li et al. (2023) fine-tune small parameter models like T5 using counterfactual contexts, SFR-RAG  Nguyen et al. (2024)is a 9B models trained with various RAG-domain datasets. However, such fine-tuning approaches require substantial computational resources and are often difficult to transfer to other models. Among non-fine-tuning approaches, one category focuses on carefully designed prompt templates. For example, One method reconstruct the context as the narrator’s statement and then inquire about the narrator’s viewpoint, encouraging the model to seek external perspectives Zhou et al. (2023). Another method employs "Role Play" interventions to alter LLMs’ decision-making styles when facing knowledge conflicts. Ying et al. (2024) However, these methods heavily rely on the model’s inherent reasoning abilities and often suffer from limited generalizability. Another category of non-fine-tuning approaches modifies the model’s decoding strategy to improve its reliance on context. CADShi et al. (2023)contrasts output probabilities with and without context to enhance contrastive decoding, amplifying the contextual probability distribution without accounting for conflicting contexts. In contrast, COIECDYuan et al. (2024) detects knowledge conflicts by measuring entropy changes during generation and dynamically adjusts the decoding strategy accordingly.
@@ 7 Conclusions
While FaithfulRAG advances error mitigation in text-based retrieval-augmented generation, its current scope is limited to textual inputs and does not yet support multimodal information (e.g., images, audio, or structured data). Extending the framework to incorporate multimodal input would enable a more comprehensive assessment of how models navigate and resolve conflicts between heterogeneous knowledge sources. Given that real-world information is often multimodal, such an extension could improve FaithfulRAG’s ability to detect and reconcile inconsistencies that arise from modality-specific biases or divergent contextual signals.
In addition, multimodal integration would enhance the applicability of the framework to domains where information synthesis from multiple sources is crucial, such as medical diagnosis (where textual reports, imaging data and patient history must be jointly considered), autonomous systems (which rely on the integration of visual, auditory, and textual signals for decision making), and interactive AI (where understanding user intent often involves processing speech, gestures, and textual input). Addressing these challenges would not only improve FaithfulRAG’s robustness but also contribute to broader advancements in multimodal retrieval-augmented generation, cross-modal reasoning, and trustworthy AI.
@@ A.2 What are advantages of FaithfulRAG?
Robustness in Non-Conflict Scenarios. A good faithful model should improve performance in knowledge-conflict scenarios while maintaining desirable performance in non-conflict scenarios. While most existing methods suffer in non-conflict settings (e.g., KRE’s 13.8% drop on SQuAD-golden), our FaithfulRAG achieves the best performance on both SQuAD-golden and MuSiQue-golden. It gets 96.6% on SQuAD-golden (+1.4% over origin) and 85.3 on MuSiQue-golden (+1.3%), proving its robustness and capability to avoid overfitting to either LLM’s parametric knowledge or incorrect contextual information.
@@
Despite recent advances in retrieval mechanisms, a persistent challenge remains: a significant gap exists between the quality of the retrieved context and the factual accuracy of the generated output. While the context retrieved by the model may be highly relevant, the generative models often struggle to faithfully incorporate this context, resulting in generated text that may diverge from factual correctness. Recent empirical evidence reveals the critical gap between retrieval quality and generation faithfulness. The generation stage determines how this context is interpreted and integrated with the model’s parametric knowledge. Existing methods often fail at this stage due to (i) Over-confidence: many RAG systems rigidly prioritize the retrieved context by suppressing parametric knowledge, leading to blind acceptance of flawed or conflicting information. (ii) Logical incoherence: Without critical reasoning, models struggle to resolve contradictions, resulting in inconsistent or unsafe outputs. By targeting generation, in this paper, we address the root cause of unfaithfulness and guide the LLM to dynamically reconcile knowledge conflicts.
@@ C.4 Implementation Details of Ablation Study
Finally, we conduct an ablation study on Reasoning stage. We substituted the structured reasoning with a naive Chain-of-Thought (CoT) approach, which does not provide explicit structured guidance for the LLM’s reasoning. This modification makes the model rely solely on its implicit inference capabilities. The prompt is as follows: